Budgeted Costs Number of Employees
S1 $4,101,000 87
S2 3,110,000 58
S3 2,865,000 37
P1 204
P2 306

Service department costs are allocated by the direct method. The number of employees is used as the allocation base for all service department costs.

Collapse question part

Allocate service department costs to production departments.

Service Service dept. Cost Allocated to
Department Costs P1 P2
S1 $4,101,000 $

$

S2 3,110,000 $

S3 2,865,000 $
Answer #1

Number of employees in department P1 = 204

Number of employees in department P2 = 306

Total employees in department P1, and P2 = Number of employees in department P1 + Number of employees in department P2

= 204+306

= 510

Percentage of employees of department P1 = Number of employees in department P1 / Total employees in department P1, and P2

= 204/510

= 40%

Percentage of employees of department P2 = Number of employees in department P2 / Total employees in department P1, and P2

= 306/510

= 60%

Service Service dept. Cost Allocated to
Department Costs P1 P2
S1 $4,101,000

4,101,000 x 40% = 1,640,400

4,101,000 x 60% = 2,460,600

S2 3,110,000

3,110,000 x 40% = 1,244,000

3,110,000 x 60% = 1,866,000

S3 2,865,000 2,865,000 x 40% = 1,146,000 2,865,000 x 60% =1,719,000

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks